lanaa:

positive a^3+b^3=(a+b)(a^2-ab+b^2) negative a^3-b^3=(a-b)(a^2+ab+b^2) 4.) x^3+1000=0 (factor the cubic) (5 and 6 factor the polynomial) 5.) 8x^3+64=0 6.)216x^3+27=0

lanaa:

@adrianluvvsyouu2

adrianluvvsyouu2:

Whats the question?

lanaa:

4-6

lanaa:

theyre numbered its factoring polynomials

adrianluvvsyouu2:

The fourth one is basically the formula I used earlier a^3 - b^3 = (a-b)(a^2 - ab + b^2) a^3 + b^3 = (a+b)(a^2 - ab + b^2) <-- Try using this one Basically you need to change 1000 into a number to the third power, said number is b, and x is a.

adrianluvvsyouu2:

Same thing with the other two, change the values into a number to the third power (cubed) So 8 is the same as 2^3.. 64 is the same as 4^3.. this way you can plug them into this formula every single time without fail

lanaa:

im confused

adrianluvvsyouu2:

@lanaa wrote:
im confused
What's confusing, let me know

lanaa:

how do i write it

adrianluvvsyouu2:

You want it in either of these forms (a+b)(a^2 - ab + b^2) or (a-b)(a^2 - ab + b^2), depending on original expression.. if the two terms have a plus in between them, you use the formula that has (a+b), if the two terms have a minus between them you use the formula that has (a-b) Let me use an example with the first problem.. We have x^3 + 1000 The first thing I want to do is get it into the form a^3 + b^3, so that the formula works Basically, I already have x^3, now I just need to change 1000 into a number to the third power. 1000 is the same as 10^3, so I want to rewrite this as x^3 + 10^3, now it's in the same form as a^3 + b^3 where x = a and 10 = b Then I can plug this into the larger formula (a+b)(a^2 - ab + b^2) notice how I used the formula with the (a+b)? This is because the original expression had a plus between the two terms So now I can say, for every a in that formula, I just need to put an x For every b in that formula, I just need to put a 10... (x + 10)(x^2 - x * 10 + 10^2) Simplified to (x+10)(x^2 - 10x + 100)

lanaa:

thank you can u help me with the other 2

adrianluvvsyouu2:

@lanaa wrote:
thank you can u help me with the other 2
I want you to use the same process, you won't be able to pass your quiz if you can't do it yourself. Let me know if you get stuck and where you got stuck

lanaa:

its not even a quiz its ok. ill find someone else.

adrianluvvsyouu2:

@lanaa wrote:
its not even a quiz its ok. ill find someone else.
You just want quick answers then instead of learning the material? How will you ever understand if you refuse to learn... it's against the rules to just provide you with the complete answer regardless.

Spectrum:

Hey! I'm learning this in Algebra 2 right now, I can definitely help you. :3

lanaa:

thank youuuu

Spectrum:

I can help you with 5 and 6 first, I'm much more familiar with these.

Spectrum:

#5: Your equation is \[8x^3+64=0\], and you must use the formula to find the sum of cubes (since you are adding.) and the formula is: \[a^3 + b^3 = (a+b)(a^2 − ab + b^2 )\]

Spectrum:

So first we take out the GCF (Greatest Common Factor), what's the biggest number that can go into both 8 and 64? (You tell me. :p)

grimreaper124:

what they said

V4MPZSKII:

w

v4mpz:

W